OpenStudy (anonymous):

plz help....The area of the batter's box on a major league baseball field is 24ft. The length of the batter's box is 2ft more than the width . Find the length and width of a rectangular batter's box

OpenStudy (anonymous):

area =l *w l = w+2 24 = (w+2) *w \[24 = w ^{2} +2w\] \[w ^{2} +2w-24 =0\] \[w ^{2}+6w-4w-24 =0\] \[w(w+6)-4(w+6) =0\] (w-4)(w+6)=0 therefore w=4 or w=-6 but dimension cannot be -ve so w=4 ft so can you find length?

OpenStudy (anonymous):

I am stuck how do l find the length

OpenStudy (anonymous):

I've already told you: l = w+2

OpenStudy (anonymous):

o o my bad l get it hold on

OpenStudy (anonymous):

the length is 6 ft

OpenStudy (anonymous):

hey l just google this site not sure how it works tho so is this a open or closed discussion

OpenStudy (anonymous):

its open

OpenStudy (anonymous):

ok so l can l ask you another question:)

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I'll be back in 10min

OpenStudy (anonymous):

how can i determine the area of the figure using the generic units defined below ...........so the small box to the right says = 1 square unit